OpenStudy (anonymous):

If f(x) = -x + 1 and g(x) = 4x - 2, what is (f-g)(x)?

OpenStudy (anonymous):

(f-g)(x)=f(x)-g(x)=-5x+3

OpenStudy (anonymous):

yes, thank you